Charms

Relic of Nature’s Aid

Cost: 1 mote Duration: Instant Type: Reflexive Minimum Martial Arts: 2 Minimum Essence: 2 Prerequisite Charms: None

The character takes a single natural, formerlyliving object, no heavier than her fist, which must have an ’edge’ of some kind, and imbues it with essence. This hardens and sharpens it into a lethal weapon. For the rest of the scene, she may wield it as a weapon equivalent to a throwing dagger.

White and Crimson Circle

Cost: 3 motes Duration: Instant Type: Simple Minimum Martial Arts: 3 Minimum Essence: 2 Prerequisite Charms:

Relic of Nature’s Aid

The character throws a daisy, or other such small circular flower. This counts as an attack with Relic of Nature’s Aid. However, if this attack does at least one point of damage, the flower has penetrated the target’s body. If the flower does this, it then explodes, the petals flying outwards at high speed in circle of tiny, vicious shards. This causes additional levels of lethal damage equal to the character’s Essence. The victim can roll Stamina and each success subtracts one.

Death Flies on Butterfly Wings

Cost: 5 motes Duration: Instant Type: Simple Minimum Martial Arts: 4 Minimum Essence: 2 Prerequisite Charms:

White and Crimson Circle

The character takes out a dead butterfly, kisses it, and sends it flying towards her target, animating it with the power of essence. This counts as an attack with Relic of Nature’s Aid. However, the butterfly seeks out its target’s anima, swirling and fluttering around any defence as it attacks. The enemy must subtract a number of successes from any defence roll equal to their own Essence, as the more Essence they wield, the more skillfully the butterfly attacks. The enemy’s soak is also reduced by their Essence against this attack.

Calming Lotus Eyes

Cost: None Duration: Permanent Type: Special Minimum Martial Arts: 3 Minimum Essence: 2 Prerequisite Charms: None

The character has mastered the art of appearing reassuringly innocent and harmless, even when innocence is the last thing on her mind. She may double her initiative on the first round of combat, because her opponents do not suspect her of such ferocity. This may, depending on circumstances, also have an effect in social or larcenous situations.

Folding Orchid Technique

Cost: 5 motes Duration: Instant Type: Supplemental Minimum Martial Arts: 5 Minimum Essence: 2 Prerequisite Charms:

Calming Lotus Eyes

Striking at crucial essence chakras on her target’s body, the character closes the victim’s mind to the outside world like a flower closes as dusk falls. A target struck by the Folding Orchid Technique loses their ability to sense their surroundings; the world becomes dim and blurry, sounds become distant and pain recedes to a dull ache.

If this attack is successful, it does not cause the loss of health levels. Instead, it removes the opponent’s sensory abilities. They gain a dice pool penalty equal to the character’s Essence on all rolls which require sensory abilities, unless they can avoid this by making a Wits + Essence roll at a difficulty equal to the character’s Essence. This lasts for a number of rounds equal to the number of damage successes. Repeated applications of this charm cause the duration to stack. Extras hit by this charm are automatically rendered helpless for a scene.

Indigo Butterfly Form

Cost: 6 motes

Duration: One scene

Type: Simple

Minimum Martial Arts: 4

Minimum Essence: 2 Prerequisite Charms:

White and Crimson Circle, Folding Orchid Technique

Weaving a complex kata, the character begins to move as gracefully as a butterfly. Her limbs and weapons slice through air and flesh alike like steel wings as her anima begins to gently flow in a hypnotic fluttering pattern behind her. She attacks from unexpected angles, leaping around defense and striking at vulnerable positions before dancing back out of range.

For the duration of the scene, the character may add her Martial Arts in dice to all attack and damage rolls, and applicable ’natural’ objects touched by her are under the effects of Relic of Nature’s Aid. A character can only use one Martial Art form type charm at one time.

Feather In The Wind Technique

Cost: 5 motes Duration: Instant Type: Reflexive Minimum Martial Arts: 4 Minimum Essence: 2 Prerequisite Charms:

Indigo Butterfly Form

The character leaps up into the air and is swept away as if by a strong wind, her body floating as if as light as a feather. This charm doubles her jumping distance horizontally, and four times as high vertically. The charm remains active as long as the character is in the air, and allows her to fall up to several hundred yards without the risk of damage. In addition, this can be used to augment a dodge : add the character’s Martial Arts to the dodge roll.

Tulip-in-Winter Stratagem

Cost: 7 motes Duration: Special Type: Reflexive Minimum Martial Arts: 5 Minimum Essence: 3 Prerequisite Charms:

Feather In The Wind Technique

When she is successfully hit in combat, the character may choose to pretend the enemy has struck a killing blow. The character rolls her Martial Arts and subtracts the successes from the damage done by the attack, although her wounds appear to be serious. She falls to the floor and appears to all observers to be dead. Her skin grows cold and her lips turn blue. If an attempt is made to determine if she is dead or alive, a Perception + Medicine roll of difficulty 4 is required to find out the truth.

In this dormant state, the character does not regain essence or health levels. Her metabolism has reduced to a tiny trickle; she is immune to poison, drowning, hunger, thirst or exposure. Her Bashing and Lethal soak is increased by her Martial Arts.

The character hears and feels nothing for the remainder of the scene after the charm is activated and may not awaken during this time. From the end of the scene, she becomes vaguely, instinctively aware of her surroundings although none of her senses are active. She can detect any presences nearby or imminent danger. She may remain dormant for up to her Martial Arts + Endurance + Essence hours, at which point she awakens quietly.

Falling Wasp Strike

Cost: 2 motes per die Duration: Instant Type: Simple Minimum Martial Arts: 5 Minimum Essence: 2 Prerequisite Charms:

Indigo Butterfly Form

Leaping high into the air, the character draws her essence around her before plunging down towards her target, energy blazing around her foot or weapon as she falls in a lethal violet arc. This attack sacrifices accuracy, as the character’s descent gives the enemy more time to avoid the attack, but woe to any who fail to get out of the way...

When the character declares this attack she decides how many motes to spend on it. There is no limit on this number. For every 2 motes spent, she loses 1 success on her attack roll but the base damage of the attack is increased by 3, if she hits. This attack can only be used with bare-handed or melee weapon based attacks, not ranged attacks. It requires at least 1 yard of headroom per mote spent upon it, so may be unsuitable for use in cramped locations. It negates any possible “perfect attack” effects; if the character gets 0 successes on her attack roll with this charm, she always misses.

If the character botches this attack, she takes one level of unsoakable bashing damage, as she slams her hand or foot into the ground very very hard.

Perfect Ending

Cost: 3 motes plus 1 per die Duration: Instant Type: Supplementary Minimum Martial Arts: 5 Minimum Essence: 3 Prerequisite Charms:

Falling Wasp Strike

In the moment of ending, all things are perfect, as they finally reach the moment they existed so long to meet. By imbuing a once-living object with a kiss of Essence, a character can cause it to become, for a second, perfection realised, greater than it ever was in life. This charm is typically used with Relic of Nature’s Aid or its related techniques, although it could be used in other situations, such as a Performance. It causes a dead leaf, flower or other such object to suddenly become a glowing ( typically purple or deep blue ), ethereal icon of its former living self, breathtakingly and tragically beautiful. After just a few seconds under the influence of this charm, the target object is totally subsumed back into the essence flow of the world and gone forever

This charm can always be used with Relic of Nature’s Aid and related charms, without the use of a combo. It cannot be used with barehanded attacks, and will destroy any weapon it is used with. This charm adds dice to the attack roll equal to the number of additional motes spent upon it. In addition, up to [half the number of motes spent] extra successes count twice for determining the damage of the attack. The maximum number of additional motes that can be spent on this charm is the character’s [Dexterity + Martial Arts].

Violet Autumn Storm Style

Cost: 10 motes, 1 Willpower, 1 health level Duration: One scene Type: Simple Minimum Martial Arts: 5 Minimum Essence: 4 Prerequisite Charms: Perfect Ending, Tulip-in-Winter Stratagem

Moving with the violence and grace of a feather spiraling in a tornado, the character is surrounded by a roiling purple and red aura filled with circling violet butterflies, which whips around her opponent as she assails them with endless fury like a relentless storm.

As long as the character spends all the actions on attacking one individual, she may split her dice pool into two attacks at -0/-0, three at -1/-1/-1, etc. up to a maximum equal to her (Essence x 2) attacks. Each of these attacks may be used to perform a Relic of Nature’s Aid, White and Crimson Circle or Death Flies on Butterfly Wings attack without requiring the use of a Combo, although the full cost of these charms must be paid, and all the attacks must be of the same type if this is done.

In addition, the storm automatically causes [the character’s Essence] - [her opponents Essence] + 1 health levels of Lethal damage to her chosen opponent at the end of each round. This effect has a range of 10 yards.